OpenStudy (anonymous):

Prove that w^49+w^101+w^150=0,where w is a complex cube root of unity

hartnn (hartnn):

now this is simple

hartnn (hartnn):

w^3 =1 did u know this ?

OpenStudy (anonymous):

REALLY!

OpenStudy (anonymous):

-1?

OpenStudy (anonymous):

yes i know w^3 =1 ,w+w^2+1=0 and etc

hartnn (hartnn):

so w^49 = ?

hartnn (hartnn):

w^49 = w^48 . w = (w^3)^16 . w = ??

OpenStudy (anonymous):

w

hartnn (hartnn):

similarly try to find w^101

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

w^101=> w^100.w => (w^3)^33.w.w=>w^2

hartnn (hartnn):

good :) w^150 = ?

OpenStudy (anonymous):

w^150 => (w^3)^50 => 1

hartnn (hartnn):

so?

hartnn (hartnn):

w^49+w^101+w^150=?

OpenStudy (anonymous):

w^2+w+1=0

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

0

hartnn (hartnn):

welcome :)